OpenStudy (anonymous):

A pyramid with height 15 is separated into two pieces by a plane parallel to the base and 6 units above it. What is the positive difference in the volumes of these two pieces if the volume of the original pyramid is 250 cubic units?

OpenStudy (anonymous):

I feel like I'm answering my own questions now, but I'm getting 196. How about the rest of you?

OpenStudy (anonymous):

Vol = 1/3 b h. Right? Original height is 15 and small one's height is 6. And base sides should be in the same proportion (2/5) So I'd say the volume of the small one should be (2/5)^3 times volume of the large one. Is that right?

OpenStudy (phi):

the top is similar to the whole, with height 15-6=9 so 9^2/ 15^2 = x/250 to find the vol of the top we can then find the bottom, and the difference

OpenStudy (anonymous):

Wait, don't you have to divide 250 by 15 first to get the area of a base?

OpenStudy (anonymous):

Oh and multiply by 3?

OpenStudy (phi):

amend that to 9^3/ 15^3 = x/250

OpenStudy (anonymous):

I think so.

OpenStudy (anonymous):

I am getting 54 for that and once you subtract it, you get 196?

OpenStudy (anonymous):

That's what I get too.

OpenStudy (anonymous):

But I just checked the answer key, and it says that it's 142

OpenStudy (phi):

the top is 54 , the bottom is 250- 54= 196 the difference between the two is 196-54= 142

OpenStudy (anonymous):

Wait, why would you subtract twice?

OpenStudy (anonymous):

Wouldn't 54 be the volume of one piece and 250 the other?

OpenStudy (phi):

the question is asking for the difference between the vol of the top piece and the bottom piece.

OpenStudy (phi):

250 is the entire pyramid

OpenStudy (anonymous):

Oh! I just misread it...now I understand...thank you!